# Loss-Guided Auxiliary Agents for Overcoming Mode Collapse in GFlowNets

### v1

@@ Abstract
Although Generative Flow Networks (GFlowNets) are designed to capture multiple modes of a reward function, they often suffer from mode collapse in practice, getting trapped in early discovered modes and requiring prolonged training to find diverse solutions. Existing exploration techniques may rely on heuristic novelty signals. We propose Loss-Guided GFlowNets (LGGFN), a novel approach where an auxiliary GFlowNet’s exploration is directly driven by the main GFlowNet’s training loss. By prioritizing trajectories where the main model exhibits high loss, LGGFN focuses sampling on poorly understood regions of the state space. This targeted exploration significantly accelerates the discovery of diverse, high-reward samples. Empirically, across various benchmarks including grid environments, structured sequence generation, and Bayesian structure learning, LGGFN consistently enhances exploration efficiency and sample diversity compared to baselines. For instance, on a challenging sequence generation task, it discovered over 40 times more unique valid modes while simultaneously reducing the exploration error metric by approximately 99%.
[…]
7899
1 Introduction
@@
Tasks such as material design and drug discovery are fundamental to advancing essential scientific fields such as medicine [23], energy, engineering and technology [13]. However, these tasks face the challenge of exploring vast combinatorial design spaces, ranging from $10^{60}$ possible drug-like molecules to potentially more for functional materials, making exhaustive search infeasible. Traditionally, drug and material discovery relied heavily on time-consuming and costly experimental screening and expert-guided design. [54, 29, 12].
@@
To overcome these limitations, machine learning has emerged as a key tool for accelerating discovery by providing proxy models [45, 9, 26] or methods to guide experiments [17, 18]. Moreover, it enabled better exploration of large combinatorial spaces efficiently by leveraging their underlying structure. Reinforcement Learning (RL) [50] is the most popular learning framework for this purpose, as it enables agents to iteratively interact with an environment to discover the single candidate that maximizes a reward, making it well-suited for tasks where the goal is to identify one optimal solution. However, in a real-world setting, a high-scoring candidate may still pose undesirable side effects or be synthetically inaccessible. Therefore, instead of identifying a single optimal solution, it is often more practical to generate a diverse set of high-quality candidates, increasing the chances that at least one will be viable in practice.
@@
Generative Flow Networks (GFlowNets) [4] aim at learning a forward policy $P_{F}(a\mid s)$, encoding choices of actions $a$ at partial states $s$, that sequentially constructs an object $x$ with probability proportional to a given reward $R(x)>0$. Unlike RL, which prioritizes finding the single best candidate, GFlowNets focus on generating a diverse set of high-reward samples. GFlowNets have found great success in many areas of applications, such as causal discovery [36, 14], material discovery [11, 37], drug discovery [47, 41], biological sequence design [24] and editing [16], large language model improvement [51, 21, 22], diffusion models improvement [57], scheduling [58] and combinatorial optimization problems [56] . In addition, alternative theories that extend GFlowNets to environments that are continuous [30], stochastic [40], and adversarial [25], which further expand the scope of possible applications such as an alternative to denoising diffusion modes [46].
@@
Even if GFlowNets and Reinforcement Learning have two different objectives, they share their need for exploration to achieve their goal. In sparse reward settings, an RL agent that doesn’t explore will not be able to improve its “understanding” of the environment and will converge to a suboptimal policy. In the same manner, even though they were specifically designed to capture multiple modes of a reward function, GFlowNets training strategies that lack exploration might lead the agent to get stuck in a region and make the underlying distribution incomplete. This behavior is more visible in large sparse-reward environments [35] where the agent might need to go through expansive regions with no or minimal reward.
@@
Previous studies have aimed to enhance GFlowNet training by investigating credit assignment [34, 32, 38]. However, akin to RL, the effectiveness of GFlowNets relies on the trajectories used to explore high-reward regions. Due to this similarity, prior work on GFlowNets [39, 44, 31, 27, 33] has drawn insights from the RL literature [8, 7, 53] to develop new exploration strategies. In particular, Bengio et al. [5] briefly cite the idea of using a second GFlowNet trained mostly to match a different reward function that is high when the losses observed by the main GFlowNet are large. Building on this idea, Madan et al. [33] suggested to implement this idea by using novelty-based intrinsic rewards, whereas Kim et al. [28] directly uses the loss value of a trajectory.
@@
In this work, we introduce a new general training strategy for GFlowNets that can be used with any training objective. Instead of using an artificial novelty-based intrinsic reward to judge how informative a trajectory is, we directly use the loss value. To avoid ruining the objective probability distribution, we use an auxiliary agent that will sample never-seen trajectories. In Sections 3 and 2, we introduce GFlowNets and review existing exploration strategies used in their training. Section 4 outlines the motivation behind our approach, details of the proposed method, and the advantages of leveraging the loss signal over traditional novelty-based intrinsic rewards. Finally, in Section 5, we show that our method matches or exceeds the previous state-of-the-art across diverse experimental settings, achieving improvements ranging from 10% in one task to 99% in another where prior methods failed entirely.
2 Background
@@
We consider a directed acyclic graph $({\mathcal{S}},{\mathcal{A}})$. We refer to the nodes ${\mathcal{S}}$ as states and to the edges ${\mathbb{A}}$ as actions. If $(s,s^{\prime})\in{\mathcal{A}}$, we say that $s$ is a parent of $s^{\prime}$ and that $s^{\prime}$ is a child of $s$. There is a unique state $s_{0}$ that has no parents that we name source state and a unique state with no child $s_{f}$ that we name sink state. We refer to the parents of the sink state as terminating states ${\mathcal{X}}=\{s\in{\mathcal{S}}\mid(s,s_{f})\in{\mathcal{A}})\}$. A complete trajectory $\tau=(s_{0},s_{1},...,s_{n},s_{n+1}=s_{f})$ is a sequence of states that start with the source state and ends with the sink state.
@@
We also consider a reward function $R:{\mathcal{X}}\rightarrow{\mathbb{R}}$ such as $\forall x\in{\mathcal{X}},R(x)>0$. By abuse of notation, we can also consider that the reward for non terminating states is zero $\forall s\in{\mathcal{S}}\setminus{\mathcal{X}},R(s)=0$ The goal of the GFlowNet is to learn to sample terminating states proportionally to the reward function $P_{T}(x)\propto R(x)$. To do this, GFlowNets start from the initial state (for example, empty molecule in drug discovery setting, initial coordinates in hyper-grid environment) and iteratively sample actions to move to the following states.
@@
Formally, we consider a non-negative flow function $F:{\mathcal{A}}\rightarrow{\mathbb{R}}$. The goal of the GFlowNet framework is to learn such a flow function that satisfies the following flow matching (FM) and reward matching constraints:
$\forall s^{\prime}\in{\mathcal{S}}\setminus\{s_{0},s_{f}\},\quad\sum\limits_{(s,s^{\prime})\in{\mathcal{A}}}F(s\rightarrow s^{\prime})$ $=\sum\limits_{(s^{\prime},s^{\prime\prime})\in{\mathcal{A}}}F(s^{\prime}\rightarrow s^{\prime\prime})$ (1)
$\forall s^{\prime}\in{\mathcal{X}},\quad F(s^{\prime}\rightarrow s_{f})$ $=R(s^{\prime})$ (2)
@@
$\forall s\in{\mathcal{S}},\quad F(s)=\sum\limits_{(s,s^{\prime})\in{\mathcal{A}}}F(s\rightarrow s^{\prime})$ (3)
@@
$P_{F}(s^{\prime}\mid s)=\frac{F(s\rightarrow s^{\prime})}{F(s)}$ (4)
$P_{B}(s\mid s^{\prime})=\frac{F(s\rightarrow s^{\prime})}{F(s^{\prime})}$ (5)
The FM constraint ensures that for every non-terminal state $s$, the total incoming flow equals the total outgoing flow. Inspired by the FM objective, detailed balance (DB) condition is a stricter version that enforces equality of flows between individual state pairs as shown in Equation 6 below:
$\forall(s,s^{\prime})\in{\mathcal{A}}\text{ and }s^{\prime}\neq s_{f},\quad F(s)P_{F}(s^{\prime}\mid s)$ $=F(s^{\prime})P_{B}(s\mid s^{\prime})$ (6)
$\forall s\in{\mathcal{X}},\quad F(s)P_{F}(s_{f}\mid s)$ $=R(s)$ (7)
In practice, DB implies FM since satisfying pairwise flow equality naturally results in balanced total inflows and outflows at each state. Therefore, the FM and DB constraints are mathematically equivalent when applied over every edge in both directions and all states [6].
Although initial work on GFlowNets focused primarily on FM and DB objectives, in this paper, we will use the trajectory balance (TB) objective to train GFlowNets. Introduced in Malkin et al. [34], the trajectory-decomposable loss only parametrizes the forward and backward policy, and add a learnable scalar $Z_{\theta}$ that represents the total reward:
${\mathcal{L}}_{TB}(\tau,\theta)=\left(\log\left(\frac{Z_{\theta}\prod_{t=1}^{n+1}P^{\theta}_{F}(s_{t}\mid s_{t-1})}{R(s_{n})\prod_{t=1}^{n}P^{\theta}_{B}(s_{t-1}\mid s_{t})}\right)\right)^{2}.$ (8)
Unlike the FM objective, TB requires only trajectory-level rewards rather than fine-grained supervision at each state or transition, making it more scalable to complex environments. Similarly, compared to the DB objective, TB avoids the need to compute flows for all intermediate states explicitly and is easier to optimize in practice due to its trajectory-wise formulation. Additionally, TB naturally accommodates stochastic policies and enables learning from sampled trajectories without requiring full state space enumeration.
3 Related Work
Exploration in Reinforcement Learning: Exploration is a critical aspect of reinforcement learning (RL). Traditional methods, such as $\varepsilon$-greedy [50], optimism-based upper confidence bounds (UCB) [1], and Thompson sampling [52], often struggle in complex environments. Approaches based on counts and pseudo-counts [3] reduce the focus on frequently visited states, but they fail to scale in high-dimensional settings. Other intrinsic motivation (IM) methods leverage indirect novelty signals as intrinsic rewards. For instance, Random Network Distillation [7] uses the discrepancy between a target and a trainable neural network to capture novelty. instead of absolute novelty, NovelD [59] uses novelty difference to give large intrinsic reward at the boundary between explored and unexplored regions. Go-Explore [15] takes a different approach by storing promising states to facilitate easy returns to these regions. Other approaches also used an auxiliary agent to help the main learner improve [2, 20, 49].
Training GFlowNets: GFlowNets were introduced by [4], employing the flow-matching objective to reformulate the flow-matching and reward-matching constraints as a loss function. To improve credit assignment and stabilize training, various alternative objectives have been proposed. The detailed-balance objective [5] operates at the level of individual transitions, while the trajectory-balance objective [34] evaluates entire trajectories. Additionally, the SubTB($\lambda$)-objective [32] enables learning from incomplete trajectories, boosting training efficiency. Recent work suggest that weighting the DB objective based on the number of terminal descendants improves the training process [48].
Exploration in GFlowNets: The success of GFlowNet training largely depends on the quality of the sampled trajectories. If training is focused on trajectories that consistently converge to a restricted region, sampling diversity can suffer. To address this, [39] introduced intrinsic rewards based on novelty methods [7, 42] to foster exploration of diverse states. Expanding on this idea, [33] proposed a secondary GFlowNet, driven by intrinsic rewards, to help the primary GFlowNet better approximate the target distribution. DGFN [31] is another approach that involves an auxiliary GFlowNet that is responsible for sampling trajectories, similar to DDQN [53]. In an effort to enhance exploration, [44] adapted Thompson sampling for GFlowNets. Another approach, suggested by [27], involves refining training trajectories using a local search algorithm before presenting them to the GFlowNet for learning.
4 Methodology
@@ 4.1 Motivation
To understand why GFlowNets can fail to learn the desired distribution, we consider a simple illustrative environment, depicted in Figure 1. The state space is defined as ${\mathcal{S}}={s_{0},s_{1},\dots,s_{N}}\cup{s_{f}}$, and the action space is ${\mathcal{A}}={\texttt{advance},\texttt{exit}}$. For every $i\leq N$, the action exit leads from $s_{i}$ to the terminal state $s_{f}$, and for every $i<N$, the action advance leads from $s_{i}$ to $s_{i+1}$. The reward function assigns high values to the extremal states, with $R(s_{0})=R(s_{N})\gg R(s_{i})$ for all $i\in{1,\dots,N-1}$. We use a tabular GFlowNet where each state $s$ is mapped to a parameter $\theta_{s}$ such that $P_{F}(\texttt{exit}|s)=\sigma(\theta_{s})$ (where $\sigma$ is the sigmoid function) and a parameter $Z$ represents the total reward learned.
Figure 2: Plot of the evolution of $P_{F}(\texttt{exit}\mid s_{0})$ for different algorithms. The chain has 100 states and uses the reward setup $R(s_{0})=R(s_{N})=101$ and $R(s_{i})=1$. At convergence, $P_{F}(\texttt{exit}\mid s_{0})=\frac{101}{300}\approx 0.337$.
Despite its simplicity—the environment has a linear, non-branching DAG structure, a single non-terminal action, and a binary reward pattern—training a GFlowNet in an on-policy manner on this environment fails to converge within a reasonable number of iterations. Intuitively, for the GFlowNet to approximate the correct distribution, it must discover and propagate the high reward associated with $s_{N}$. However, assuming the initial policy is a uniform action policy where each action is chosen with probability 0.5, the probability of reaching $s_{N}$ is $0.5^{N}$. This probability diminishes exponentially with $N$, and as training progresses, the policy increasingly prioritizes the more easily discovered high-reward state $s_{0}$, further reducing the chance of ever reaching $s_{N}$. As a result, $s_{N}$ is typically discovered far later than expected (approximately 190,000 trajectories in the case of a chain of length 100, as indicated by a noticeable inflection point in the green curve in Figure 2), which significantly slows down learning despite the simplicity of the environment.
@@
4.2 Algorithm
In the example illustrated in Figure 1, once the model has learned to terminate trajectories at $s_{0}$, such trajectories become dominant, making the discovery of the final state slower. Figure 2 shows that even after 200,000 trajectories, training using an on-policy method (TB) has yet to converge for a 100-sized chain. Moreover, standard exploration techniques such as tempering or $\varepsilon$-greedy exploration prove insufficient in overcoming this limitation.
@@ 4.2 Algorithm
where $R_{\text{main}}$ is the original reward used to train the main GFlowNet, and $\mathcal{L}_{\text{main}}$ denotes the loss associated with a state or trajectory under the main GFlowNet. This auxiliary reward prioritizes trajectories that the main GFlowNet has not yet learned (i.e., those with high loss), while reducing focus on already well-learned regions (i.e., low-loss trajectories). Even if we mainly focus on trajectory-based objective, the same idea holds for transition, state or sub-trajectory based objective. As a result, this encourages broader exploration and mitigates mode collapse.
To leverage the auxiliary GFlowNet, we train the main agent using a mixture of on-policy trajectories and trajectories sampled by the auxiliary GFlowNet. This strategy helps the main agent retain knowledge of previously discovered modes while also exposing it to regions where its performance remains weak. To ensure training stability, we include $R_{\text{main}}$ in the auxiliary reward to provide a consistent structural signal, thereby counteracting the inherent fluctuations of the loss term. The coefficient $\lambda$ balances the influence of the loss and the reward, ensuring that the overall scale of $R_{\text{aux}}$ remains comparable to that of $R_{\text{main}}$ and avoids destabilizing the learning process. The training procedure is outlined in detail in Algorithm 1.
@@
3:     Sample trajectories $\tau_{\text{aux}}\sim F_{\text{aux}}$
@@
5:     Compute loss $\mathcal{L}(\tau_{\text{aux}},F_{\text{main}},R)$
6:     Compute auxiliary reward: $R_{\text{aux}}=R+\lambda\mathcal{L}(\tau_{\text{aux}},F_{\text{main}},R)$
7:     Update $F_{\text{aux}}$ using $\mathcal{L}(\tau_{\text{aux}},F_{\text{aux}},R_{\text{aux}})$
@@
9:     Update $F_{\text{main}}$ using $\mathcal{L}(\tau,F_{\text{main}},R)$
@@
4.3 Benefits of loss-guided auxiliary GFlowNet
@@
Using the loss of the main GFlowNet as a guide for the auxiliary GFlowNet allows to leverage the generalization induced by the use of Neural Networks. During training, neural networks learn to recognize patterns and generalize them to previously unseen states. For instance, as discussed in Section 5.2, the model may infer that when a sequence is incomplete but valid, the optimal next action leading to high-reward states is consistently appending a 00. Due to this generalization, some trajectories or transitions—despite never being explicitly encountered—can still produce a low loss. Leveraging the loss to guide exploration takes advantage of this property.
@@
In contrast, relying solely on novelty-based intrinsic rewards might drive the model toward areas that are technically new but were already implicitly understood through the network’s generalization. Using the loss is also cheaper computationally, since it will be computed anyway to backpropagate on the main agent, whereas intrinsic rewards ask for a separate set of calculations, and can be expensive depending on which novelty-based algorithm is used.
@@
While similar in spirit, the reward formulation proposed in Kim et al. [28] differs in two aspects. First, it is asymmetric, placing greater emphasis on trajectories where the estimated forward reward $ZP_{F}(\tau)$ is lower than the expected reward $R(x_{\tau})P_{B}(\tau\mid x_{\tau})$. Second, it combines the primary and auxiliary objectives multiplicatively, using the formulation $\log{\mathcal{R}}_{\text{aux}}=\alpha\log{\mathcal{R}}+\log{\mathcal{L}}$. However, our experiments indicate that this added complexity does not lead to improved performance. In fact, our simplified variant outperforms it by a small margin.
5 Experiments
@@
In the experiments section, we compare three GFlowNets training procedures with the trajectory balance objective: on-policy (that we simply refer to as TB), Siblings Augmented GFlowNet [33] (SAGFN), and our loss-guided auxiliary GFlowNet (LGGFN). We first validate the ability of both SAGFN and LGGFN to discover all modes in the grid, then we show that LGGFN outperforms SAGFN in settings where the rewards have a strong structure, as we predicted in 4.3.
5.1 Hypergrid

### v2

@@ Abstract
Although Generative Flow Networks (GFlowNets) are designed to capture multiple modes of a reward function, they often suffer from mode collapse in practice, getting trapped in early-discovered modes and requiring prolonged training to find diverse solutions. Existing exploration techniques often rely on heuristic novelty signals. We propose Loss-Guided GFlowNets (LGGFN), a novel approach where an auxiliary GFlowNet’s exploration is directly driven by the main GFlowNet’s training loss. By prioritizing trajectories where the main model exhibits high loss, LGGFN focuses sampling on poorly understood regions of the state space. This targeted exploration significantly accelerates the discovery of diverse, high-reward samples. Empirically, across diverse benchmarks including grid environments, structured sequence generation, Bayesian structure learning, and biological sequence design, LGGFN consistently outperforms baselines in exploration efficiency and sample diversity. For instance, on a challenging sequence generation task, it discovered over 40 times more unique valid modes while simultaneously reducing the exploration error metric by approximately 99%.
Introduction
@@
Tasks such as material design and drug discovery are fundamental to advancing essential scientific fields such as medicine (Hughes2011DrugDiscovery), energy, engineering and technology (Curtarolo2013HighThroughput). However, these tasks involve exploring vast combinatorial design spaces, ranging from $10^{60}$ possible drug-like molecules to potentially more for functional materials, making exhaustive search infeasible. Traditionally, drug and material discovery relied heavily on time-consuming and costly experimental screening and expert-guided design. (Wani1971Taxol; Klayman1985Artemisinin; Cobb2010StainlessSteel).
@@
To overcome these limitations, machine learning has emerged as a key tool for accelerating discovery by providing proxy models (Schutt2018Schnet; Chen2019GraphNetworks; Jumper2021AlphaFold) or methods to guide experiments (Hase2018Phoenics; Hase2021Gryffin). Moreover, it enables efficient exploration of large combinatorial spaces efficiently by leveraging their underlying structure. Reinforcement Learning (RL) (sutton1998reinforcement) is the most popular learning framework for this purpose, as it enables agents to iteratively interact with an environment to discover the single candidate that maximizes a reward, making it well-suited for tasks where the goal is to identify one optimal solution. However, in a real-world setting, high-scoring candidates may have undesirable side effects or be synthetically inaccessible. Therefore, generating a diverse set of high-quality candidates is often more practical than identifying a single optimal solution, increasing the chances that at least one will be viable in practice.
@@
Generative Flow Networks (GFlowNets) (bengio2021flow) aim at learning a forward policy $P_{F}(a\mid s)$, encoding choices of actions $a$ at partial states $s$, that sequentially constructs an object $x$ with probability proportional to a given reward $R(x)>0$. Unlike RL, which prioritizes finding the single best candidate, GFlowNets focus on generating a diverse set of high-reward samples. GFlowNets have found great success in many areas of applications, such as causal discovery (manta2023gflownets; deleu2022bayesian), material discovery (D4DD00020J; nguyen2023hierarchical), drug discovery (shen2024tacogfn; pandey2024gflownet), biological sequence design (jain2022biological) and editing (ghari2024gflownet), large language model improvement (Takase2024GFlowNet; ho2024proof; hu2024amortizing; younsi2025accurate), diffusion models improvement (DBLP:journals/corr/abs-2406-00633), scheduling (zhang2023robust) and combinatorial optimization problems (zhang2023let) . In addition, alternative theories that extend GFlowNets to environments that are continuous (lahlou2023theory), stochastic (pan2023stochastic), and adversarial (jiralerspong2024expected), which further expand the scope of possible applications such as an alternative to denoising diffusion modes (sendera2024improved).
@@
Even if GFlowNets and Reinforcement Learning have two different objectives, they share their need for exploration to achieve their goal. In sparse reward settings, an RL agent that doesn’t explore will not be able to improve its “understanding” of the environment and will converge to a suboptimal policy. In the same manner, even though they were specifically designed to capture multiple modes of a reward function, GFlowNets training strategies that lack exploration might lead the agent to get stuck in a region and make the underlying distribution incomplete. This behavior is more visible in large sparse-reward environments (malkin2023gflownets) where the agent might need to go through expansive regions with no or minimal reward.
@@
Previous studies have aimed to enhance GFlowNet training by investigating credit assignment (malkin2022trajectory; madan2023learning; pan2023better). However, akin to RL, the effectiveness of GFlowNets relies on the trajectories used to explore high-reward regions. Due to this similarity, prior work on GFlowNets (pan2023generative; rector2023thompson; lau2023dgfn; kim2023local; madantowards) has drawn insights from the RL literature (chapelle2011empirical; burda2018exploration; van2016deep) to develop new exploration strategies. In particular, JMLR:v24:22-0364 briefly cite the idea of using a second GFlowNet trained mostly to match a different reward function that is high when the losses observed by the main GFlowNet are large. Building on this idea, madantowards suggested to implement this idea by using novelty-based intrinsic rewards, whereas kim2025adaptive directly uses the loss value of a trajectory.
@@
In this work, we introduce a new general training strategy for GFlowNets that can be used with any training objective. Instead of using an artificial novelty-based intrinsic reward to judge how informative a trajectory is, we directly use the loss value. To avoid ruining the objective probability distribution, we use an auxiliary agent that will sample never-seen trajectories. The Background section introduces GFlowNets, while the related works section in the appendix reviews existing exploration strategies employed in their training. Our methodology section presents the motivation for our approach, provides detailed descriptions of the proposed method, and examines the benefits of utilizing loss signals compared to conventional novelty-based intrinsic rewards. The Experiments section then shows that our method matches or exceeds the previous state-of-the-art across diverse experimental settings, including hypergrid environments of varying sizes and sparsity levels, structured bit sequence generation, Bayesian network structure learning, and a novel multi-objective mRNA sequence design environment, achieving improvements ranging from 10% in Bayesian structure learning to 99% reduction in exploration error in sequence generation where prior methods failed entirely.
The method shows particular strength in sparse reward settings where traditional on-policy training completely fails, as demonstrated by our motivating example where discovering a distant mode requires navigating through exponentially many low-reward states.
Background
Readers unfamiliar with GFlowNets are encouraged to refer to the first section of the Appendix, which provides additional context and introduces some key ideas to understand this framework.
@@
We consider a directed acyclic graph $({\mathcal{S}},{\mathcal{A}})$, where nodes $\mathcal{S}$} represent states and edges $\mathcal{A}$} represent actions. If $(s,s^{\prime})\in{\mathcal{A}}$, we say that $s$ is a parent of $s^{\prime}$ and that $s^{\prime}$ is a child of $s$. There is a unique state $s_{0}$ with no parents that we name source state and a unique state $s_{f}$ with no children that we name sink state. We refer to the parents of the sink state as terminating states ${\mathcal{X}}=\{s\in{\mathcal{S}}\mid(s,s_{f})\in{\mathcal{A}})\}$. A complete trajectory $\tau=(s_{0},s_{1},...,s_{n},s_{n+1}=s_{f})$ is a sequence of states that start with the source state and ends with the sink state.
@@
We also consider a reward function $R:{\mathcal{X}}\rightarrow\mathbf{R}$ such as $\forall x\in{\mathcal{X}},R(x)>0$. By convention, we can also consider that the reward for non terminating states is zero $\forall s\in{\mathcal{S}}\setminus{\mathcal{X}},R(s)=0$. The goal of the GFlowNet is to learn to sample terminating states proportionally to the reward function $P_{T}(x)\propto R(x)$. To do this, GFlowNets start from the initial state (for example, empty molecule in drug discovery setting, initial coordinates in hyper-grid environment) and iteratively sample actions to move to the following states.
@@
Formally, we consider a non-negative flow function $F:{\mathcal{A}}\rightarrow\mathbf{R}$. The goal of the GFlowNet framework is to learn such a flow function that satisfies the following flow matching (FM) and reward matching constraints, for all $s^{\prime}\neq s_{0},s_{f}$ and for all $x\in{\mathcal{X}}$, respectively:
$\sum_{(s,s^{\prime})\in{\mathcal{A}}}F(s\rightarrow s^{\prime})=\sum_{(s^{\prime},s^{\prime\prime})\in{\mathcal{A}}}F(s^{\prime}\rightarrow s^{\prime\prime}),$ (1)
$F(x\rightarrow s_{f})=R(x).$ (2)
@@
$\forall s\in{\mathcal{S}},\quad F(s)=\sum\limits_{(s,s^{\prime})\in{\mathcal{A}}}F(s\rightarrow s^{\prime}).$ (3)
@@
$P_{F}(s^{\prime}\mid s)=\frac{F(s\rightarrow s^{\prime})}{F(s)},\ \ P_{B}(s\mid s^{\prime})=\frac{F(s\rightarrow s^{\prime})}{F(s^{\prime})}.$ (4)
In this paper, we use the trajectory balance (TB) objective to train GFlowNets. Introduced in malkin2022trajectory, the trajectory-decomposable loss only parametrizes the forward $P^{\theta}_{F}$ and backward $P^{\theta}_{B}$ policies, and adds a learnable scalar $Z_{\theta}$ that represents the unknown partition function. The loss for each trajectory $\tau$:
${\mathcal{L}}_{TB}(\tau;\theta)=\left(\log\left(\frac{Z_{\theta}\prod_{t=1}^{n+1}P^{\theta}_{F}(s_{t}\mid s_{t-1})}{R(s_{n})\prod_{t=1}^{n}P^{\theta}_{B}(s_{t-1}\mid s_{t})}\right)\right)^{2},$
is minimized by stochastic gradient descent, using trajectories sampled from a given behavior policy. Common choices of the behavior policy include the learned policy $P^{\theta}_{F}$ and artificially modified versions of it (bengio2023gflownet). In this work, we tackle the problem of behavior policy design, which we discuss next.
Methodology
@@
To understand why GFlowNets can fail to learn the desired distribution, we consider a simple illustrative environment, depicted in Figure 1. The state space is defined as ${\mathcal{S}}=\{s_{0},s_{1},\dots,s_{N}\}\cup{s_{f}}$, and the action space is ${\mathcal{A}}=\{\texttt{advance},\texttt{exit}\}$. For every $i\leq N$, the action exit leads from $s_{i}$ to the terminal state $s_{f}$, and for every $i<N$, the action advance leads from $s_{i}$ to $s_{i+1}$. The reward function assigns high values to the extremal states, with $R(s_{0})=R(s_{N})\gg R(s_{i})$ for all $i\in{1,\dots,N-1}$. We use a tabular GFlowNet where each state $s$ is mapped to a parameter $\theta_{s}$ such that $P_{F}(\texttt{exit}|s)=\sigma(\theta_{s})$ (where $\sigma$ is the sigmoid function) and a parameter $Z$ that represents the total reward learned.
Figure 2: Plot of the evolution of $P_{F}(\texttt{exit}\mid s_{0})$ for different algorithms as a function of number of sampled trajectories. The chain has 100 states and uses the reward setup $R(s_{0})=R(s_{N})=101$ and $R(s_{i})=1$. At convergence, $P_{F}(\texttt{exit}\mid s_{0})=\frac{101}{300}\approx 0.337$.
Despite its simplicity (linear, non-branching DAG structure with a single non-terminal action and binary reward pattern), on-policy GFlowNet training fails to converge within a reasonable number of iterations. Intuitively, for the GFlowNet to approximate the correct distribution, it must discover and propagate the high reward associated with $s_{N}$. However, assuming the initial policy is a uniform action policy where each action is chosen with probability 0.5, the probability of reaching $s_{N}$ is $0.5^{N}$. This probability diminishes exponentially with $N$, and as training progresses, the policy increasingly prioritizes the more easily discovered high-reward state $s_{0}$, further reducing the chance of ever reaching $s_{N}$. As a result, $s_{N}$ is typically discovered far later than expected (approximately 190,000 trajectories in the case of a chain of length 100, as indicated by a noticeable inflection point in the green curve in Figure 2), which significantly slows down learning despite the simplicity of the environment.
@@
Algorithm
In the example illustrated in Figure 1, once the model has learned to terminate trajectories at $s_{0}$, such trajectories become dominant on policy, making the discovery of the final state slower. Figure 2 shows that even after 200,000 trajectories, training using an on-policy method (TB) has yet to converge for a 100-sized chain. Moreover, standard exploration techniques such as tempering or $\varepsilon$-greedy exploration prove insufficient in overcoming this limitation.
@@
where $R_{\text{main}}$ is the original reward used to train the main GFlowNet, and $\mathcal{L}_{\text{main}}$ denotes the loss associated with a state or trajectory under the main GFlowNet. This auxiliary reward prioritizes trajectories that the main GFlowNet has not yet learned (i.e., those with high loss), while reducing focus on already well-learned regions (i.e., low-loss trajectories). While we mainly focus on trajectory-based objectives in this paper, the same idea holds for transition, state or sub-trajectory -based objective. As a result, this encourages broader exploration and mitigates mode collapse.
To leverage the auxiliary GFlowNet, we train the main agent using a mixture of on-policy trajectories and trajectories sampled by the auxiliary GFlowNet. This strategy helps the main agent retain knowledge of previously discovered modes while also exposing it to regions where its performance remains weak. To mitigate the training instability of the auxiliary gflownet that is due to a moving reward, we include $R_{\text{main}}$ in the auxiliary reward to provide a consistent structural signal, thereby counteracting the inherent fluctuations of the loss term. The coefficient $\lambda$ balances the influence of the loss and the reward, ensuring that the overall scale of $R_{\text{aux}}$ remains comparable to that of $R_{\text{main}}$ and avoids destabilizing the learning process. The training procedure is outlined in detail in Algorithm 1.
@@
3:  Sample trajectories $\tau_{\text{aux}}\sim F_{\text{aux}}$, with reward $R$
@@
5:  Compute main loss $\mathcal{L}(\tau_{\text{aux}};F_{\text{main}},R)$
6:  Compute auxiliary reward: $R_{\text{aux}}=R+\lambda\mathcal{L}(\tau_{\text{aux}};F_{\text{main}},R)$
7:  Update $F_{\text{aux}}$ using $\mathcal{L}(\tau_{\text{aux}};F_{\text{aux}},R_{\text{aux}})$
@@
9:  Update $F_{\text{main}}$ using $\mathcal{L}(\tau;F_{\text{main}},R)$
@@
Benefits of loss-guided auxiliary GFlowNet
@@
Using the loss of the main GFlowNet as a guide for the auxiliary GFlowNet allows to leverage the generalization induced by the use of Neural Networks. During training, neural networks learn to recognize patterns and generalize them to previously unseen states. For instance, as discussed in the BitSequence subsection in the experiments, the model may infer that when a sequence is incomplete but valid, the optimal next action leading to high-reward states is consistently appending a 0 (an open parentheses). Due to this generalization, some trajectories or transitions, despite never being explicitly encountered, can still produce a low loss. Leveraging the loss to guide exploration takes advantage of this property.
@@
In contrast, relying solely on novelty-based intrinsic rewards might drive the model toward areas that are technically new but were already implicitly understood through the network’s generalization. Using the loss is also cheaper computationally, since it will be computed anyway to train the main agent, whereas intrinsic rewards ask for a separate set of calculations, and can be expensive depending on which novelty-based algorithm is used.
@@
While similar in spirit, the reward formulation proposed in kim2025adaptive differs in two aspects. First, it is asymmetric, placing greater emphasis on trajectories where the estimated forward reward $ZP_{F}(\tau)$ is lower than the expected reward $R(x_{\tau})P_{B}(\tau\mid x_{\tau})$. Second, it combines the primary and auxiliary objectives multiplicatively, using the formulation $\log{\mathcal{R}}_{\text{aux}}=\alpha\log{\mathcal{R}}+\log{\mathcal{L}}$. However, our experiments indicate that this added complexity does not lead to improved performance. In fact, our simplified variant outperforms it by a small margin.
A short theoretical discussion of why loss-guided sampling prevents partial-support stationary points and how it implicitly induces a curriculum over trajectories is provided in the Appendix.
Experiments
@@
In the experiments section, we compare three GFlowNets training procedures with the trajectory balance objective: on-policy (that we simply refer to as TB), Siblings Augmented GFlowNet (SAGFN; madantowards), and our loss-guided auxiliary GFlowNet (LGGFN). We first validate the ability of both SAGFN and LGGFN to discover all modes in the hyper-grid environemnt introduced in bengio2021flow, then we show that LGGFN outperforms SAGFN in settings where the rewards have a strong structure: a bit-sequence environment similar to (malkin2022trajectory)’s, a causal structure learning environment (deleu2022bayesian), and a structured biological sequence design setting..
Hypergrid